OpenStudy (anonymous):

Find the zeros of the function f(x)=x^2-1/x+4

OpenStudy (abb0t):

Is your function \(x^2 - \frac{1}{x}+4\) or \(\large \frac{ x^2-1}{ x+4}\)

OpenStudy (anonymous):

f(x)=(x^2-1)/(x+4)

OpenStudy (anonymous):

Second option you posted.

OpenStudy (abb0t):

Ok, well first, set it equal to zero.

OpenStudy (abb0t):

You have \( \huge \frac{x^2-1}{x+4}=0\)

OpenStudy (abb0t):

next, multiply both sides by \(\large x+4\)

OpenStudy (abb0t):

You will be left with \(\large x^2-1=0\) which is simply the difference \(\large (x+1)(x-1) = 0\) now you will have two solutions to solve which you can easily do yoruself: \(\large (x+1) = 0\) and \(\large (x-1)=0\) and your zero's are: x =

OpenStudy (abb0t):

The zeros of a function are the x coordinates of the x intercepts of the graph of f.

OpenStudy (anonymous):

OK, So x=-1,1?

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

Oh, You are awesome. Thank you!! :)